# GraphCogent: Overcoming LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

### v1

@@ D.2 Framework Adaptability across Different LLMs
Our experiments primarily focus on the Llama3.1-8B, but the framework’s modular design ensures compatibility with both open-source and closed-source LLMs, as confirmed by cross-LLM validation results. We select four tasks within Graph4real on middle-scale: Shortest Path, Cycle Detection, Max Flow, and Link Prediction. Table 13 reveals our framework’s adaptability. For open-source models like GLM4-9B-chat and Qwen2.5-7B-coder-Instruct, our framework significantly enhances their reasoning capabilities especially for complex tasks requiring multi-step graph operations, improving performance from near-zero accuracy in tasks such as Shortest Path and Max Flow to high accuracy. Second, based on the framework’s instruction without training, massive-scale LLMs like GPT-4o and DeepSeek-R1 achieve high performance with in-toolset tasks and out-toolset, demonstrating the universal applicability of our framework design.
Appendix E Edit distance and Cost for GEC
E.1 Edit distance
Real-world scenarios present graphs in multiple text representations including adjacency lists, symbolic notations, and linguistic descriptions. This variability makes rule-based preprocessing approaches such as regular expressions impractical for standardization. Mirroring how human external sensory systems process multimodal information, our Sensory Module transforms these varied graph inputs into standardized adjacency list representations.
The transformation process requires rigorous quality assessment due to the substantial topological information contained in real-world graphs and the cascading effects of transformation errors on downstream reasoning. We employ edge Edit distance as shown in algorithm 1, quantify the minimum number of edge additions and deletions required to transform one graph into another to evaluate difference between two graphs.
E.2 Cost for Graph Efficiency Coefficient
The selection of token count as the primary cost metric in GEC computation is motivated by three considerations. First, token-based pricing serves as the universal billing standard across LLM platforms (e.g., GPT, Claude, and DeepSeek series), making it the most feasible cost indicator for comparative studies. Second, while temporal cost was initially considered, our experiments with 100 graphs (300 edges each) across four sampling granularities revealed less than 5% variance in processing time. This temporal consistency emerges because the output lengths remain relatively stable for graph transformations. Crucially, the actual computational expense manifests in input token caused by LLMs’ input instruction (i.e., the instruction guides the LLM to transform graph information into a fixed adjacency list format). We therefore adopt normalized token counts as they precisely capture this trade-off: finer sampling granularity necessitates more frequent LLM interactions, each accumulating input instruction tokens, while coarser granularity reduces interaction frequency at the risk of topological distortion.
@@ E.2 Cost for Graph Efficiency Coefficient
1:
2: Source graph $G=(V,E)$
3: Target graph $H=(V,E^{\prime})$ with identical node set $V$
4:
5: Edit distance $d_{\text{edit}}(G,H)$ between graphs $G$ and $H$
6:Initialize empty edge sets:
@@
12:Compute edge differences:
13:for each edge $e\in E_{H}$ do
14:  if $e\notin E_{G}$ then
15:   $\mathcal{A}\leftarrow\mathcal{A}\cup\{e\}$
16:  end if
17:end for
18:for each edge $e\in E_{G}$ do
19:  if $e\notin E_{H}$ then
20:   $\mathcal{R}\leftarrow\mathcal{R}\cup\{e\}$
21:  end if
22:end for
23:Calculate final distance:
[…]
25:return $d_{\text{edit}}(G,H)$
Appendix F Graph Reasoning Task Definition
We design a total number of 21 tasks and categorize them into three classes as following:
  • Graph Structural Querying Tasks: Focus on fundamental graph properties, such as edge existence or node count, testing basic graph structural comprehension.
  • Graph Algorithmic Reasoning Tasks: Require reasoning based on classical algorithms, such as shortest path computation and maximum flow analysis, evaluating algorithmic proficiency.
  • Graph Predictive Modeling Tasks: Involve neural network-based predictions, such as node classification and link prediction, assessing predictive modeling capabilities.
Table 14: Domains and corresponding scenarios.
Domain Scenarios
Social
Information diffusion analysis
Community detection and recommendation systems
Fraudulent account detection
Influence maximization algorithms
Social network dynamics analysis
Web
Web crawler efficiency optimization
Search engine ranking optimization
Web structural integrity diagnosis
Topical community discovery
DDoS attack mitigation
Transportation
Travel route planning
Logistics delivery optimization
Urban emergency response planning
Public transit network scheduling
Shared mobility platforms
Citation
Scholarly influence tracking
Interdisciplinary research identification
Seminalpaper discovery
Literature retrieval ranking optimization
Research frontier identification
For task generation, we collect five real-world scenarios as shown in Table 14 for each of the four domains (e.g., travel planning and logistics optimization for transportation) and craft 20 prompt templates (a template example is shown in Fig. 14) to construct contextually grounded questions. Using DeepSeek-R1 DeepSeek-AI [2025], we generate graph reasoning tasks based on templates with specific scenarios, ensuring both diversity and relevance. To validate task quality, we employ a dual evaluation approach: DeepSeek-R1 for automated assessment and human for manual review. This process produces a dataset with 4200 questions distributed across the 21 tasks.
F.1 Graph Structural Querying Tasks
The Graph Structural Querying tasks in Table 15 represent fundamental operations that are universally applicable across all four domains (Web, Social, Transportation, and Citation) in our benchmark. These tasks focus on direct lookup operations with constant time complexity, avoiding complex algorithmic requirements to measure basic graph comprehension abilities. Using designed task templates, we generate domain-specific versions of each Structural Querying task that maintain the same core operation while incorporating appropriate context from each domain. The five tasks include Edge Count, Node Count, Degree Count, Edge Existence, and Node Existence. All these tasks are classified as in-toolset, with each task having a corresponding pre-built tool implementation that performs direct lookup operations on the graph data structure.
Table 15: Details of Graph Structural Querying Tasks.
Task Description Tool Algorithm Time Complexity
Edge Count Count the total number of edges in a given graph. Direct Lookup $O(1)$
Node Count Count the total number of nodes in a given graph. Direct Lookup $O(1)$
Degree Count Count the number of edges connected to a specific node in a given graph. Direct Lookup $O(1)$
Edge Existence Determine whether a specific edge exists between two nodes in a given graph. Direct Lookup $O(1)$
Node Existence Determine whether a specific node exists in a given graph. Direct Lookup $O(1)$
F.2 Graph Algorithmic Reasoning Tasks
The in-toolset Graph Algorithmic Reasoning Tasks in Table 16 extend beyond basic graph comprehension to assess multi-step reasoning capabilities. These tasks employ classical graph algorithms with varying time complexities, from linear to cubic, to evaluate different aspects of structural understanding. Tasks contain Cycle Detection, Triangle Count, Path Existence, and Shortest Path. Mirroring Structural Querying tasks, these algorithmic challenges are implemented as pre-defined tools and adapted cross-domains template for each scenario.
Table 16: Details of in-toolset Graph Algorithmic Reasoning Tasks.
Task Description Tool Algorithm Time Complexity
Cycle Detection Determine whether there exists any cycle in a given graph. Depth-First Search $O(|V|+|E|)$
Triangle Count Count the total number of triangles in a given graph. Node Iterator $O(|V|\cdot d_{\text{max}}^{2})$
Path Existence Determine whether a specific path exists between two nodes in a given graph. Depth-First Search $O(|V|+|E|)$
Shortest Path Determine the minimum distance between two nodes in a given graph. Dijkstra $O({|V|}^{2}+|E|)$
Table 17: Details of out-toolset Graph Algorithmic Reasoning Tasks.
Task Description Algorithm Time Complexity
Maximum Flow Compute maximum traffic flow from source to sink node. Exclusive for Transportation. Edmonds-Karp $O(|V|\cdot|E|^{2})$
Diameter Calculation Find the longest shortest path in the given graph. Exclusive for Transportation. Floyd-Warshall $O(|V|^{3})$
Max Core Calculation Identify the largest k-core subgraph. Exclusive for Social. Core Decomposition $O(|E|)$
Connected Components Count the number of all connected subgraphs. Exclusive for Social. Breadth-First Search $O(|V|+|E|)$
Common Neighbors Count shared connections between two nodes. Exclusive for Web. Set Intersection $O(\min(d(u),d(v)))$
PageRank Calculation Compute node PageRank based on link structure. Exclusive for Web. Power Iteration $O(k|E|)$
Reference Match Calculate the number of commonly cited papers between two papers. Exclusive for Citation. Set Intersection $O(|A|+|B|)$
Clustering Coefficient Calculate local clustering coefficient of nodes. Exclusive for Citation. Triangle Counting $O(|V|\langle k^{2}\rangle)$
The out-toolset Graph Algorithmic Reasoning Tasks present domain-specific challenges that require specialized algorithmic solutions beyond the toolset. These tasks include Maximum Flow and Diameter Calculation for Transportation, Max Core Calculation and Connected Components for Social, Common Neighbors and PageRank Calculation for Web, and Reference Match and Clustering Coefficient for Citation. Each task employs distinct algorithms with varying computational complexities tailored to their specific domain requirements, evaluating LLMs’ task-specific model generation capabilities.
F.3 Graph Predictive Modeling Tasks
The Graph Predictive Modeling Tasks further evaluate LLMs’ reasoning abilities on graphs by introducing prediction-based challenges that differ from the deterministic results of Structural Querying tasks and Algorithmic Reasoning tasks. These tasks specifically assess LLMs’ ability to leverage neural networks for predictive modeling, with each task explicitly assigned a distinct neural architecture to facilitate standardized evaluation. For Traffic Prediction from PeMS datasets Shuman et al. [2013], we preprocess and store time-series data in Numpy arrays. Similarly, Node Classification tasks from Cora Craven et al. [1998] utilize pre-computed text embeddings stored in PyG format. Unlike the traffic flow prediction task and node classification task which are derived from PeMS and Cora datasets with complete task construction and evaluation schemes, Link Prediction (constructed from SNAP Social Circles McAuley and Leskovec [2012] and Google Web Graph McAuley and Leskovec [2012]) requires first constructing both the graph and the task itself. The Link Prediction tasks employ a specially constructed dataset generated through original graph sampling with added random noise, where the task is specifically defined as evaluating whether two unconnected nodes may have latent relationships based on existing edge patterns. These predictive tasks include Traffic Prediction for Transportation, Link Prediction for both Social and Web domains, and Node Classification for Citation, each implemented with domain-specific neural architectures to assess LLMs’ predictive reasoning capabilities on graph-structured data.
Table 18: Details of Graph Predictive Modeling Tasks.
Task Description Algorithm Time Complexity
Traffic Prediction Predict future traffic flow patterns. Exclusive for Transportation. LSTM $O(T\cdot n^{2})$
Link Prediction Predict missing connections between users. Exclusive for Social. GCN $O(L\cdot|E|\cdot d^{2})$
Link Prediction Predict potential hyperlinks between pages. Exclusive for Web. GraphSAGE $O(\prod_{l=1}^{L}s_{l}\cdot d^{2})$
Node Classification Classify papers into research topics. Exclusive for Citation. GAT $O(|V|\cdot d^{2}+|E|\cdot d)$
Appendix G Limitations
Although GraphCogent outperforms other baselines, it still has several limitations. First, while the framework performs well on static graphs, it does not natively support dynamic graphs with evolving topologies or edge weights. The Buffer Module assumes static data, requiring full reconstruction for updates, which introduces computational overhead. Second, the Model Agent’s reasoning capabilities are currently limited to pre-trained models (e.g., GCN, LSTM) and deterministic tasks. Supporting more advanced models may require integrating more powerful foundation models (such as DeepSeek-V3 or GPT-4o) or additional training. Moreover, the framework struggles with NP-hard problems at larger scales due to computational complexity, and its current performance evaluation primarily relies on accuracy metrics, which may not fully assess tasks requiring both numerical results and path validity (e.g., shortest path or max flow problems). Additionally, while the framework standardizes diverse graph representations, highly specialized formats (e.g., molecular graphs) may require additional domain-specific parsing. These limitations suggest directions for future work, including enhanced support for dynamic graphs, broader model generalization, and more comprehensive evaluation metrics.
Appendix H Error Analysis
In this section, we analyze the errors of GraphCogent (Llama3.1-8B based) within the Graph4real benchmark for both in-toolset tasks and out-toolset tasks.
H.1 In-toolset tasks
For in-toolset tasks, we selected Edge Count, Triangle Count, Cycle Detection, and Path Existence, analyzing error patterns across 500 test instances at the middle scale (100 nodes). The errors were categorized into the following two types:
  • Edge Error: Structural deviations of origin graph and transformed graph caused by different edges during transformation process.
  • Tool Error: Incorrect tool selection or parameterization in the Execution Module
Table 19: Error distribution across in-toolset tasks.
Task Type Edge Error Tool Error Accuracy
Edge Count 19 2 97.8%
Triangle Count 18 1 98.8%
Cycle Detection 21 5 99.0%
Path Existence 19 13 97.4%
The results in Table 19 demonstrate that GraphCogent achieves high accuracy across in-toolset tasks, but performance varies by task type. For edge-dependent tasks (e.g., Edge Count, Triangle Count), errors primarily stem from Edge Errors, structural discrepancies introduced during graph transformation. These tasks rely heavily on precise edge information, making them sensitive to minor deviations in the adjacency list conversion. In contrast, topology-independent tasks (e.g., Cycle Detection, Path Existence) show a reversed pattern. While Edge Errors may occur, their impact is mitigated because task outcomes often do not depend on specific edge subsets. Therefore, Tool Errors dominate the main error types for these tasks.
H.2 Out-toolset tasks
For out-toolset tasks, we select Max Flow on Transportation and Link Prediction on Social, analyzing error patterns across 500 test instances at the middle-scale graph (100 nodes). We classify errors into three categories:
  • Discrimination Error: Occurs when the Reasoning Agent fails to identify the task as out-toolset task, leading to inappropriate tool calls instead of activating the Model Agent.
  • Execution Error: Arises when the task-specific model generated by Model Agent contains logical flaws or implementation issues that prevent successful execution.
  • Format Error: Applies when the task-specific model produces correct results but returns them in non-compliant formats.
Table 20: Error distribution across out-toolset tasks.
Task Type Discrimination Error Execution Error Format Error Accuracy
Max Flow 35 6 11 89.6%
Link Prediction 6 58 10 85.2%
The results are reported in Table 20. For Max Flow, discrimination errors dominate, primarily due to the Reasoning Agent’s failure to discriminate task boundaries when problems exceed the toolset’s coverage. In such cases, the Reasoning Agent incorrectly attempts tool-based solutions despite the task requiring model generation. Format errors are secondary, often manifesting as mismatched outputs (e.g., returning paths instead of the max flow value). In contrast, Link Prediction is primarily affected by execution errors, which stem from internal tensor mismatches in the generated neural models (e.g., dimension inconsistencies in GCN layers or invalid adjacency matrix operations). Both discrimination and format errors occur less frequently, suggesting that the model reliably identifies the task type and produces valid outputs.
Appendix I LLM Prompts
The prompts we used for Sensory Agent are shown in Fig. 12. Prompts for Reasoning Agent are presented in Fig. 10 and Fig. 11. Fig. 13 gives the example of prompt for Model Agent. Fig. 14 gives an example of task generation.
Figure 10: Prompt for Reasoning Agent.
Figure 11: Prompt for Toolsets.
Figure 12: Prompt for Sensory Agent.
Figure 13: Prompt for Model Agent.
Figure 14: Task generation template.
Appendix J GraphCogent Workflow
Algorithm 2 presents the workflow of GraphCogent. For the four key components of the reasoning process, we provide corresponding case examples in Fig. 15, Fig. 16, Fig. 17, and Fig. 18.

### v2

@@ 5.2. Main Result
Although our experiments primarily focus on the Llama3.1-8B, the framework’s modular design ensures compatibility with both open-source and closed-source LLMs. Table 6 reveals our framework’s adaptability. For open-source models like GLM4-9B-chat and Qwen2.5-7B-coder-Instruct, our framework significantly enhances their reasoning capabilities especially for complex tasks requiring multi-step graph reasoning, improving performance from near-zero accuracy in tasks such as Shortest Path and Max Flow to high accuracy. Second, based on the framework’s instruction without training, massive-scale LLMs like GPT-4o and DeepSeek-R1 achieve high performance with in-toolset tasks and out-toolset, demonstrating the universal applicability of our framework design.
The scalability evaluation in Table 5 on larger graphs further demonstrates the effectiveness of our framework, where it maintains stable performance across various tasks. Through ablation studies in Appendix B, we also examine the contributions of each component.
6. Conclusion
In this work, we proposed GraphCogent, a novel agent framework that enhances LLMs’ graph reasoning capabilities by mitigating their working memory constraints. Inspired by human cognitive architecture, GraphCogent integrates three key modules: Sensory Module (standardizing diverse graph text representations), Buffer Module (integrating and indexing graph data), and Execution Module (combining tool-calling and tool-creation reasoning). To evaluate LLMs on real-world graph reasoning tasks, we introduce Graph4real, a benchmark featuring large-scale graphs for four real-world domains. Experiments show that GraphCogent achieves 20% higher than existing state-of-the-art agent-based method while reducing over 30% token usage.
Appendix A Additional Experimental Settings
Corresponding LLM versions. All corresponding versions of the LLMs are presented in Table 7.
Table 7. Baseline methods and corresponding models.
LLM Type Method Base Model
Text-based Two-shot prompt (Wei et al., 2022) GPT-4o-2024-08-06 (OpenAI, 2023)
Two-shot prompt (Wei et al., 2022) Claude-3.7-sonnet-20250219 (Anthropic, 2024)
Two-shot prompt (Wei et al., 2022) DeepSeek-R1 (DeepSeek-AI, 2025)
Tool-based Graph-Toolformer (Zhang, 2023) Llama3.1-8B-Instruct (Touvron et al., 2023)
GraphTool-Instruction (Wang et al., 2025) Llama3.1-8B-Instruct (Touvron et al., 2023)
Function Calling (OpenAI, 2023) GPT-4o-2024-08-06 (OpenAI, 2023)
Agent-based GraphTeam (Li et al., 2024) GPT-4o-mini-2024-07-18 (OpenAI, 2023)
GCoder (Zhang et al., 2024) Llama3.1-8B-Instruct (Touvron et al., 2023)
GraphCogent Llama3.1-8B-Instruct (Touvron et al., 2023)
[…]
Appendix B Ablation Study
We conduct ablation studies to evaluate each module and its key components through the following research questions:
B.1. Ablation Studies on Sensory Module
RQ1: How does the Graph Verifier component enhance transformation reliability and ensure global consistency?
As shown in Fig. 3 even based on GEC optimization, the inherent uncertainty in LLM outputs can still introduce parsing inaccuracies during graph transformation. Table 8 further demonstrates the critical role of the Graph Verifier in ensuring reliable graph transformation and global consistency.
We randomly sample 100 graphs (each containing 100 nodes with 300 edges) from PeMS dataset, partitioning each graph into 50-edges subsets with different graph text representations (i.e., adjacency list, symbolic notation, and linguistic description). We evaluate edit distance during transformation and assess post-merge global consistency using two criteria: (i) Global Edit Distance (GED), the edit distance between the original graph $G$ and the merged graph $\hat{G}$; and (ii) Boundary–Edge Recall (BER), the recall of edges whose endpoints lie in different windows or touch a window boundary. The results show that introducing the Graph Verifier reduces Edit distance by approximately 90%, while maintaining operational efficiency with fewer than 10 verification triggers per transformation. The results show that introducing the Graph Verifier reduces structural errors by approximately 90%, while maintaining operational efficiency with fewer than 10 verification triggers. This retry frequency (occurring in approximately 600 LLM calls) remains acceptable while significantly reducing edit distance.
Table 8. Impact of Graph Verifier on transformation reliability and global consistency. Edit Distance is computed by summing the edit distances of all windows within each graph.
Representation Type Edit Dist. (Before) Edit Dist. (After) GED BER (%) Verifier Triggers
Adjacency List $44.8\pm 7.6$ $5.2\pm 1.1$ $0.8\pm 0.2$ $99.8\pm 0.06$ $6.2\pm 1.0$
Symbolic Notation $57.3\pm 8.9$ $5.9\pm 1.2$ $1.1\pm 0.4$ $99.7\pm 0.08$ $9.0\pm 1.3$
Linguistic Description $65.9\pm 9.1$ $5.1\pm 1.0$ $1.0\pm 0.3$ $99.8\pm 0.07$ $9.1\pm 1.2$
B.2. Ablation Studies on Buffer Module
We select four tasks within Graph4real on middle-scale. The experimental results in Table 9 show that removing the Buffer Module causes substantial performance decreases across all tasks. This result evidences the Buffer Module’s critical role in addressing LLMs’ working memory constraints for graph reasoning. Without the Buffer Module, when the graph scale exceeds the capacity of LLMs’ working memory, the LLMs exhibit substantial information loss regarding graph topology.
Table 9. Performance impact of Buffer Module removal.
Variants Shortest Path Cycle Detection Max Flow Link Prediction
GraphCogent 98.2 99.0 89.6 85.2
GraphCogent w/o Buffer 68.4 69.2 53.8 54.2
B.3. Ablation Studies on Execution Module
RQ3: Does Thinking Policy Initialization (TPI) improve Execution Agent’s tool selection?
We select four tasks within Graph4real on middle-scale. Three versions of the Execution Agent are evaluated: the original untuned agent (Llama3.1-8B), agent fine-tuned using the Naive SFT method, and agent fine-tuned with our Thinking Policy Initialization (TPI). The Naive SFT approach involves training Llama3.1-8B by directly mapping problem statements to their outputs, where each output consists of a brief analytical statement containing the tool selection result. As shown in Table 10, the TPI method shows consistent performance advantages across different tasks. For in-toolset tasks, TPI achieves a 5% performance improvement over the naive approach, as the explicit reasoning chains enable the agent to better understand task intent and select appropriate tools. Through analysis of the reasoning process, we observe that compared to Naive SFT’s tendency to directly output corresponding tools (i.e., simply fitting tool results), the TPI teaches the agent to comprehend task intentions and demonstrates reasonable reasoning paths, leading to correct tool selection. This advantage becomes even more significant for out-toolset tasks, where the TPI shows substantially better generalization capabilities than naive method.
Table 10. Impact of Thinking Policy Initialization on Execution Agent performance.
Llama3.1-8B (Base) 78.4 79.2 35.8 36.2
Llama3.1-8B w/ SFT 89.3 90.2 43.8 44.2
Llama3.1-8B w/ TPI 98.2 99.0 70.2 65.0
However, the TPI still has limitations. Through analyzing the reasoning results for out-toolset tasks like Max Flow and Link Prediction, we identify a tendency of the LLM to forcibly associate tools. For instance, although no single tool can solve the Max Flow problem, unlike the Naive SFT-tuned LLM, which randomly generates a tool or calls a non-existent tool, the TPI-tuned LLM attempts to combine tools to derive a solution. This preference actually contradicts practical decision-making. In this task, we would prefer the agent to directly generate the task-specific tool of Max Flow through Tool Creator rather than repeatedly interact with tools of Execution Agent. This observation motivates our implementation of CMPO alignment to refine the Execution Agent’s preferences.
RQ4: Why employ Capability-Margin Preference Optimization (CMPO) for tool capability discrimination?
Building upon the findings from B.3, the core objective of CMPO alignment is to refine the Execution Agent’s tool selection preferences for improving toolset discrimination capabilities. Although tasks are not strictly binary (i.e., solvable either by tools or not), the LLM requires a decision-making mechanism to optimize resource efficiency. The goal is to minimize resource consumption, single tool calls can significantly reduce the LLM’s interaction tokens. The experimental results in Table 11 demonstrate that CMPO not only preserves Execution Agent’s in-toolset selection performance but also substantially enhances out-toolset discrimination capability.
Table 11. Impact of CMPO on Execution Agent.
Execution Agent w/o CMPO 97.2 97.8 70.2 64.8
Execution Agent w/ CMPO 98.2 99.0 89.7 85.2
RQ5: How does Tool Creation Fine-tuning improve Tool Creator’s generation capability?
We select four out-toolset tasks within Graph4real on middle-scale. Three versions of the Tool Creator are evaluated: the original untuned agent (Llama3.1-8B), agent fine-tuned using Code-Enhanced SFT method, and agent fine-tuned using Tool Creation Fine-tuning method. Table 12 reports the accuracy and executable percentage. The improvements in both accuracy and executability metrics confirm the effectiveness of Tool Creation Fine-tuning. Compared to Code-Enhanced SFT, especially for mainstream open-source LLMs, it is more challenging to enable LLMs to master the reasoning capabilities for complex tasks through code training. The complexity of the tasks and the limitations of LLM capabilities hinder their generalization performance on complex tasks. In contrast, Tool Creation Fine-tuning effectively enhances accuracy and executability by allowing the agent to focus more on generating core task-specific tools.
Table 12. Impact of Tool Creation Fine-tuning.
Without SFT With Code SFT With Tool SFT
Task Acc.(%) Execut.(%) Acc.(%) Execut.(%) Acc.(%) Execut.(%)
Max Flow 38.2 58.4 72.2 81.2 89.6 96.3
Common Neighbors 42.6 63.1 70.6 78.3 88.8 95.8
Link Prediction 39.8 61.6 58.6 65.2 85.2 88.4
Node Classification 41.2 59.8 60.8 64.7 89.6 89.9
RQ6: Can GraphCogent deal with multi-step tool planning tasks?
To further probe our framework’s generalization to multi-step tool planning with interpretable reasoning paths, we introduce three classic multi-step reasoning tasks in Web scenarios: (i) Crawl-trap detection, which first detects redirect loops and plans a route; (ii) k-core filtered ranking, which computes PageRank on the k-core subgraph; and (iii) Capacity-aware pathing, which gates path computation with a capacity threshold before finding a shortest path. Each workflow is instantiated on middle-scale Web subgraphs. We compare four system variants: Naive Llama3.1-8B, TPI-only, CMPO-only, and GraphCogent (TPI+CMPO). We report two planning metrics: PlanEM, the percentage of cases where the predicted tool sequence exactly matches the reasoning chain; and TraceAlign-F1, the F1 between tool steps explicitly stated in the agent’s thinking trace and the actually executed steps. As shown in Table 13, GraphCogent achieves consistently higher PlanEM and TraceAlign-F1 across all workflows, indicating more reliable sequencing and better trace–execution agreement; compare with TPI-only and CMPO-only, margin calibration reduces boundary confusions.
Table 13. GraphCogent’s generalization to multi-step tool planning tasks. Trace-F1 denotes TraceAlign-F1.
Crawl-trap detection k-core filtered ranking Capacity-aware pathing
Method PlanEM (%) Trace-F1 (%) PlanEM (%) Trace-F1 (%) PlanEM (%) Trace-F1 (%)
Naive Llama3.1-8B 60.1 57.3 55.1 51.0 57.6 54.5
TPI-only 72.4 69.1 68.9 63.7 70.3 66.5
CMPO-only 83.8 80.3 79.7 75.9 82.5 78.8
GraphCogent 93.5 90.2 91.6 86.8 92.7 88.5
Appendix C Edit distance and Cost for GEC
@@
1:Source graph $G=(V,E)$; Target graph $H=(V,E^{\prime})$ with identical node set $V$
2:Edit distance $d_{\text{edit}}(G,H)$
3:Assert $V(H)=V(G)$
[…]
5:if $E_{G}=E_{H}$ then
6:  return 0
7:end if
@@
12:return $d_{\text{edit}}(G,H)$
C.1. Edit distance
Real-world scenarios present graphs in multiple text forms—adjacency lists, symbolic notations, and linguistic descriptions—making rule-based preprocessing (e.g., regular expressions) impractical for standardization. Following human external sensory processing, our Sensory Module converts these varied inputs into a standardized adjacency-list representation. Because small transformation errors can propagate downstream reasoning on topology-rich graphs, we perform rigorous quality assessment using edge edit distance (Algorithm 1), which measures the minimum edge additions and deletions needed to transform one graph into another.
C.2. Cost for Graph Efficiency Coefficient
We select token count as GEC’s cost metric for three reasons: (i) token-based pricing is the universal billing standard across LLM platforms (GPT, Claude, DeepSeek), enabling fair comparison; (ii) in tests on 100 graphs (300 edges) with four granularities, processing time varied by less than 5%, since transformation outputs are length-stable, so time is not discriminative; (iii) the dominant expense is input tokens from the fixed instruction used to standardize graphs. We therefore use normalized token counts to capture the granularity trade-off: finer sampling increases LLM calls and accumulates instruction tokens, whereas coarser sampling reduces calls at the risk of topological distortion.
Appendix D Error Analysis
In this section, we analyze the errors of GraphCogent within the Graph4real benchmark for both in-toolset and out-toolset tasks.
D.1. In-toolset tasks
We select four in-toolset tasks and analyze error patterns across 500 test instances at the middle scale (100 nodes). The errors are categorized into two types: (i) Edge Error: Structural deviations between the original graph and the transformed graph caused by inconsistent edges during the transformation. (ii) Tool Error: Incorrect tool selection or parameterization in the Execution Module.
Table 14. Error distribution across in-toolset tasks.
Task Type Edge Tool Accuracy (%)
Edge Count 19 2 97.8
Triangle Count 18 1 98.8
Cycle Detection 21 5 99.0
Path Existence 19 13 97.4
The results in Table 14 demonstrate that GraphCogent achieves high accuracy across in-toolset tasks, but performance varies by task type. For edge-dependent tasks (e.g., Edge Count, Triangle Count), errors primarily stem from Edge Errors, structural discrepancies introduced during graph transformation. These tasks rely on precise edge information, making them sensitive to minor deviations in the adjacency list conversion. In contrast, topology-independent tasks (e.g., Cycle Detection, Path Existence) show a reversed pattern. While Edge Errors may occur, their impact is mitigated because task outcomes do not depend on specific edge subsets. Therefore, Tool Errors dominate the error types for these tasks.
D.2. Out-toolset tasks
For out-toolset tasks, we select Max Flow on Transportation and Link Prediction on Social, analyzing error patterns across 500 test instances at the middle-scale graph (100 nodes). We classify errors into three categories: (i) Discrimination Error: Occurs when the Execution Agent fails to identify the task as an out-toolset task, leading to inappropriate tool calls instead of activating the Tool Creator. (ii) Execution Error: Arises when the task-specific tool generated by the Tool Creator contains logical flaws or implementation issues that prevent successful execution. (iii) Format Error: Arises when the task-specific tool produces correct results but returns them in non-compliant formats.
Table 15. Error distribution across out-toolset tasks.
Task Type Discrimination Execution Format Accuracy (%)
Max Flow 35 6 11 89.6
Link Prediction 6 58 10 85.2
The results are reported in Table 15. For Max Flow, discrimination errors dominate, primarily due to the Execution Agent’s failure to discriminate task boundaries when problems exceed the toolset’s coverage. In such cases, the Execution Agent incorrectly attempts tool-based solutions despite the task requiring tool creation. Format errors are secondary, often manifesting as mismatched outputs (e.g., returning paths instead of the max flow value). In contrast, Link Prediction is primarily affected by execution errors, which stem from internal tensor mismatches in the generated neural models (e.g., dimension inconsistencies in GCN layers or invalid adjacency matrix operations). Both discrimination and format errors occur less frequently, suggesting that the model reliably identifies the task type and produces valid outputs.
Appendix E GraphCogent Workflow
Algorithm 2 presents the workflow of GraphCogent. More detailed workflow and corresponding prompt templates are provided in our released anonymous github code.